New studies indicate that corn has significant healing powers if it is consumed daily. How will these studies affect the corn market? Explain the adjustment process to the new equilibrium.

Income per month Quantity of fish
Period 1 $750 8 lbs
Period 2 $950 5 lbs

a. Using the table above, compute the income elasticity of demand for fish.

b. Is fish a normal good or inferior good? Explain.

Price of fish Quantity of fish Price of broccoli Quantity of broccoli
Period 1 $7 8 lbs $4.50 10 lbs
Period 2 $9 5 lbs $6.25 6 lbs

c. Calculate the cross price elasticity of broccoli.

d. Are fish and broccoli complementary or substitute goods? Explain.
Answer #1

a) Income elasticity = [5-8/(5+8/2)] / [950-750/(950+750/2)] = [-3/6.5] / [200/850] = -0.4615/0.2352 = -1.96

b) Fish is an inferior good as income elasticity is negative so when income increases, the consumption of fish decreases

c) Cross price elasticity = [6-10/(6+10/2)] / [9-7/(9+7/2)] = [-4/8] / [2/8] = -0.5/0.25 = -2

d) Fish and broccoli are complimentary goods as the elasticity is negative